OpenStudy (anonymous):

in an alloy of 150 pounds of zinc and copper, there are 100 pounds of copper. How much copper must be added so that the alloy may be 10 % zinc?

OpenStudy (amistre64):

if we setup a ratio: p c z 150:100:50 we want to know if we add k pounds of copper 150+k :100+k :50; such that 50/(150+k) = .1

OpenStudy (anonymous):

i think it is 35pounds of copper that needs to be added...

OpenStudy (anonymous):

so i have to add 200# of copper

OpenStudy (amistre64):

well, 50 is 10% of 500, sooo 150+k=500

OpenStudy (amistre64):

id say at least 200 :) maybe more

OpenStudy (anonymous):

350

OpenStudy (anonymous):

ii have the answer for this

OpenStudy (amistre64):

350 sounds about right :) 450/500 = .90 50/500 = .10

OpenStudy (anonymous):

he got it

OpenStudy (anonymous):

ok that makes sense!! whew!!! I am a music teacher,, I count to 4 lol. Thanks for your help!!

OpenStudy (amistre64):

good luck ;)